OpenStudy (anonymous):

A helium tank has a volume of 2.6 L, a pressure of 3.50 * 10^4 kPa, and is in a room kept at 21 degrees Celsius. The balloons being filled each have a maximum volume of 2.55 L and maintain a temperature and pressure at 30 degrees Celsius and 115 kPa respectively. How many balloons is the tank capable of filling? Would these balloons pop if 1.5 moles of helium is added to them? I figured out that the tank can fill 320.94 balloons, but how do you solve the second question?

OpenStudy (somy):

well maximum volume for the balloon is V= 2.55L/1000= 0.00255 m3 P=115 000 Pa T= 30 Celsius +273= 303 Kelvin R= 8.31 PV=nRT n=RT/PV n= (8.31*303)/(115000*0.00255) you'll get the mole if the this mole is LESS than 1.5 moles then the balloon will pop if its MORE than 1.5 moles then nothing's gonna happen that's how i understand the question Hope that helps :D Feel free to ask if something is not clear :)